Shift report 20000812 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Aug 12 2000 - 07:41:53 EDT

  • Next message: http account: "Shift report 20000812 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Aug 12 07:41:53 2000, Shift: 00:00-08:00
    Supervisor: EJK/RW, with: B4
    Log book#4 pages 58 - 61
    
    Shift Summary:
    ------------------------------------------------------------
    1. We finally got beam at 5:10am for about 45min. Beam 
    was 55 bunches at the full energy and the intensity was 
    6300e6 in Blue and 7300e6 in yellow. After 45min. Cryo PLC 
    fault was detected and beam was stopped. They started to 
    reinject beam on 7:00am, but did not yet succeeded.
    2. Run 2201 is finished. 14395 collisions are contained.
    3. TOFW seems no data on SuperMonitor, although HV are ON.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Aug 12 2000 - 07:45:41 EDT